[(D-1)^2.(D^2+1)^2]y=sin^2x/2+e^x
1
Expert's answer
2020-10-07T18:52:43-0400

The Auxiliary equation is given by,

"[(m-1)^2(m+1)^2]=0"


The roots of the above equation are,

1,1,-1,-1

Therefore The complimentary function,

C.F.="(c_1+c_2x)e^x+(c_3+c_4x)e^{-x}" ......(1)

Particular Integral, P.I.= "\frac{sin^2x}{2(D^2-1)^2}+\frac{e^x}{(D^2-1)^2}"


= "\frac{sin^2x}{2(-1-1)^2}+\frac{xe^x}{2(D^2-1)\times 2D}"

="\frac{sin^2x}{2(-2)^2}+\frac{x^2e^x}{4(3D^2-1)}"


="\frac{sin^2x}{8}+\frac{e^x}{4(3.1-1)}"

="\frac{sin^2x}{8}+\frac{e^x}{8}"

So complete solution=C.F+P.I.

y ="(c_1+c_2x)e^x+(c_3+c_4x)e^{-x}+\frac{sin^2x+e^x}{8}"